What is the equation of the line that passes through (-3, -1) and has a slope of ? Put your answer in slope-intercept form.
Answer by Alan3354(69443)

What is the equation of the line that passes through (-3, -1) and has a slope of ? Put your answer in slope-intercept form.
----------
Use m for slope
y = ms + b is the slope-intercept form
=======================
y+1 = m*(x+3) = mx + 3m
y = mx + (3m-1)
3m-1 is the y-intercept
==============
PS Don't put a space after the comma.
(-3,-1), not (-3, -1)